There are fi ve old ATM machines at Smart Bank. All probability that an ATM breaks
down in the course of a day is 0.2. If they are all in working order at the beginning a
day, and breakdowns are independent of each other, fi nd the probability that during
the day;
i. there are no breakdowns
ii. four machines break down
iii. all machines break down
iv. at least three machines break down
1
Expert's answer
2020-04-06T13:07:51-0400

Let p be the probability of a ATM breakdown,

p=0.2


i. there are no breakdowns,

"P_1=(1-p)^5\\
P_1=(0.8)^5\\
P_1=0.32768"


ii. four machines break down

"P_2=(p^4 )\times (1- p)\\
P_2=(0.2)^4 \times (0.8)\\
P_2=0.00128"



iii. all machines break down

"P_3=p^5\\
P_3=(0.2)^5\\
P_3=0.00032"


iv. at least three machines break down

"P_4=P_2+P_3+(p^3\times(1-p)^2)\\
P_4=0.00128+0.00032+0.2^3\times0.8^2\\
P_4=0.00672"